OpenStudy (cucacula):

approximate √36.03 by using the concept of derivative. anyone?

OpenStudy (agent0smith):

Let x = 36, so y =6 from \[\large y = \sqrt x\] so \[\Large \frac{ dy }{ dx } = \frac{ 1 }{ 2 \sqrt x }\] \[\Large dy = \frac{ 1 }{ 2 \sqrt x } dx\] \[\Large \Delta y = \frac{ 1 }{ 2 \sqrt x } \Delta x\] Again, x =36 so let the Δx be 0.03. Find Δy, then you can approximate the sqrt of 36.03 by adding Δy to y (which is 6).

OpenStudy (cucacula):

how to find dx? @agent0smith

OpenStudy (agent0smith):

If x is 36, dx is 0.03, since we want to approximate sqrt 36.03, so 36.03-36 = 0.03.

OpenStudy (agent0smith):

dx just means the change in x. x changes from 36 to 36.03, and we use 36 as x since we know the sqrt of 36 is 6.

OpenStudy (cucacula):

 so it was like this? @agent0smith